OpenStudy (anonymous):

Suppose that a and b vary inversely and that b=5/3 when a = 9. Write a function that models the inverse variation and find b when a = 10.

OpenStudy (kenljw):

b = A/a A constant 5/3 = A/9 therefore A = 15 b = 15/a

OpenStudy (anonymous):

b=15/15?

OpenStudy (kenljw):

at a = 15

OpenStudy (kenljw):

b = 15/10 = 3/2

OpenStudy (anonymous):

Nevermind, I was just confused, Thank You

OpenStudy (kenljw):

The result of equation should determine choice.